OpenStudy (anonymous):

Write an exponential function y = abx for a graph that includes (2, 24) and (3, 48). y = 2(6)x y = 4(3)x y = 3(4)x y = 6(2)x

OpenStudy (anonymous):

I will give medal

OpenStudy (anonymous):

First, plug in the values of x=2 and y=24 for the first equation, and x=3 and y=48 for the second one: 24 = ab^2 48 = ab^3 divide equation 2 by equation 1: 2 = b^1 or, more simply: 2 = b now plug b into one of the equations to get "a": 24 = a*2^2 24 = a*4 6 = a So, the answer is "C" y = 6(2)^x (I think you need to specify the power of x in your answers)

OpenStudy (anonymous):

can you help me with some more please? @Hannah_Waller

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Suppose the population of a town is 9,800 and is growing 2% each year. Write an equation to model the population growth. Predict the population after 7 years. y=9,800*1.02^x; about 11,257 people y=9,800*2^x; about 11,257 people y=9,800*2^x; about 1,254,400 people y=2* 9,800^x; about 137,200 people @Hannah_Waller

OpenStudy (anonymous):

I think its the 2nd or third one...

OpenStudy (anonymous):

Suppose an investment of $6,600 doubles in value every 8 years. How much is the investment worth after 40 years? $105,600 $66,000 $528,000 $211,200 @Hannah_Waller

OpenStudy (anonymous):

The last one

OpenStudy (anonymous):

Write an exponential function y = abx for a graph that includes (–3, 16) and (–1, 4) f(x) = 3(4)x f(x) = 2(0.5)x f(x) = 0.5(2)x f(x) = 4(0.3)x

OpenStudy (anonymous):

Honestly im not completly sue on this one.... @tay30

OpenStudy (anonymous):

Oh that is ok. Suppose a laboratory has a 31 g sample of polonium-210. The half-life of polonium-210 is about 138 days. How many half-lives of polonium-210 occur in 966 days? How much polonium is in the sample 966 days later? 7; 0.24 g 7; 2,139 g 8; 0.12 g 7; 4.43 g

OpenStudy (anonymous):

i think mabye d but idk...

OpenStudy (anonymous):

Write an exponential function y = abx for a graph that includes (1, 15) and (0, 6). y = 6(2.5)x y = 2.5(6)x y = 3(5)x y = 5(3)x

OpenStudy (anonymous):

@Hannah_Waller

OpenStudy (anonymous):

im not good at these... sorry

OpenStudy (anonymous):

Ok. Thanks for the help. @Hannah_Waller